# ADAM-SINDy: An Efficient Optimization Framework for Parameterized Nonlinear Dynamical System Identification

### v2

@@ 4 Discussion
Interestingly, the ADAM-SINDy framework also facilitates interaction among the functions within the candidate library through the self-adaptation of nonlinear functions. This self-adaptive mechanism significantly enhances the model’s capacity to capture complex dynamical behaviors, thereby allowing it to adjust to the varying degrees of nonlinearity inherent in real-world systems. Such adaptability is particularly crucial when predefined candidates are lacking, as it optimizes parameters to compensate for absent functions. In contrast, the classical SINDy approach is constrained by static nonlinear parameters and a lack of interaction, increasing the risk of inaccurate system models when essential candidate functions are omitted. This adaptability parallels the advantages observed in symbolic regression techniques, further enriching the model’s applicability. However, the occurrence of self-adaptation may transform the candidate matrix prior to optimization into an “ill-conditioned” matrix following optimization. Based on our tests, ill-conditioned cases do not pose a problem for ADAM-SINDy and simply result in two identical terms (in different forms) that need to be manually added for simplification of the final equation (see Fig. 2).
@@
Concurrently, ADAM-SINDy aims to identify a parsimonious mathematical model while optimizing trainable sparsity-promoting hyperparameters, such as $\bm{\Gamma}$. A similar adaptive weighting strategy can be found in other research works, such as PINN [57] and Transformers [67]. From the sensitivity analysis in Sec. 3.7, it is evident that both trainable hyperparameters can be employed for ADAM-SINDy optimization. However, prioritizing the trainable $\bm{\Gamma}$ is recommended, as it reliably produces the correct terms with machine precision, whereas $\lambda$ may identify the correct terms but with slight deviations in their numerical values. While ADAM-SINDy offers both batch-wise and full-batch optimization, it has been observed that for simpler dynamical systems, a small randomly sampled batch of data is often sufficient to capture the necessary temporal information for accurately estimating governing parameters. However, for more complex systems, a small batch may not encompass all required information, necessitating a larger batch size that ideally approaches full-batch size. Consequently, full-batch optimization is preferred and utilized in the present work, leading to reduced uncertainty in the converged optimized parameters within the ADAM-SINDy methodology.
@@
It is worth noting that symbolic regression via neural networks has been employed to identify dynamical systems [60]. Boddupalli et al. [60] applied this approach to chemical reaction kinetics, the same system considered in this work (Sec. 3.4), resulting in an overly complex, “bloated” mathematical model when explicit stacking within the neural network is absent. However, when exponential terms are explicitly stacked with polynomials, the identified model closely approximates the ground truth equation but exhibits slight deviations in the coefficient values. Remarkably, ADAM-SINDy accurately identified the true dynamical system, capturing the dynamics without requiring additional treatments to ensure regression convergence.
@@
In real-world problems, the presence of noise in the observational data from various sources is inevitable, prompting an investigation into the efficacy of the ADAM-SINDy method. To assess its robustness against noisy data, we compared the ADAM-SINDy methodology to the classical SINDy-W method by applying Gaussian white noise of up to 50% of the standard deviation to all the numerical example datasets. However, the results from these noise level experiments were not presented in this paper, as both methods demonstrated similar performance and failed to exhibit advantages at high noise levels. Both methods performed effectively at lower noise levels, reinforcing the notion that neither method outperforms the other in the presence of significant noise. Future work will focus on further improving the ADAM-SINDy method to achieve superior performance under noisy conditions.
Furthermore, it is essential to acknowledge a notable limitation of the full batch-ADAM-SINDy methodology: its inherent computational expense due to the utilization of the ADAM optimization algorithm for the coefficient matrix. Although this approach offers significant advantages over the classical SINDy method, it incurs a considerable computational cost. This cost arises despite the fact that the ADAM-SINDy method mirrors all computational strategies of its predecessor, differing primarily in the employed optimization algorithm. Efficiency can be achieved by integrating an iterative least squares method for optimizing the coefficient matrix while utilizing the ADAM optimizer for nonlinear parameters and sparsity knobs, leading to a combined least-squares-ADAM optimization framework for system identification. However, integrating these approaches presents challenges, particularly in ensuring compatibility between the iterative least squares method and the ADAM optimizer. Achieving seamless integration requires addressing potential conflicts between sparsity constraints and nonlinear optimization objectives. Consequently, further research is essential to develop robust methodologies that can effectively merge these distinct approaches while preserving the desirable properties of each. Additionally, one can explore the efficacy of different algorithms, including variations of ADAM, to assess their performance in this context.
In the context of the ABC flow example, efforts were made to leverage the computational graph of the gradient descent algorithm by incorporating a divergence-free constraint into the loss equation. This approach demonstrated a noteworthy reduction in the total number of epochs required for optimization—approximately 40 epochs—compared to the typical requirement of around $6\times 10^{4}$ epochs. This observation highlights the potential for further enhancement of the method in future work. Thus, a promising direction for future research involves exploring differential equation-based sparsity controllers. Such advancements could facilitate the development of a more parsimonious and interpretable dynamical system model that aligns with underlying physical laws, thereby significantly improving the computational efficacy of the ADAM-SINDy framework.
@@ Acknowledgments
The ADAM-SINDy framework, developed in a PyTorch environment, will be made available on GitHub following the peer review process.

### v3

@@ 4 Discussion
Interestingly, the ADAM-SINDy framework also facilitates interaction among the functions within the candidate library through the self-adaptation of nonlinear functions. This self-adaptive mechanism significantly enhances the model’s capacity to capture complex dynamical behaviors, thereby allowing it to adjust to the varying degrees of nonlinearity inherent in real-world systems. Such adaptability is particularly crucial when predefined candidates are lacking, as it optimizes parameters to compensate for absent functions. In contrast, the classical SINDy approach is constrained by static nonlinear parameters and a lack of interaction, increasing the risk of inaccurate system models when essential candidate functions are omitted. This adaptability parallels the advantages observed in symbolic regression techniques, further enriching the model’s applicability. However, the occurrence of self-adaptation may transform the candidate matrix prior to optimization into an “ill-conditioned” matrix following optimization. Based on our tests, ill-conditioned cases do not pose a problem for ADAM-SINDy and simply result in two identical terms (in different forms) that need to be manually added for simplification of the final equation.
@@
Concurrently, ADAM-SINDy seeks to identify a parsimonious mathematical model while optimizing trainable sparsity-promoting hyperparameters, such as $\bm{\Gamma}$, through two distinct objective formulations: joint-minimization and min-max. While the min-max adaptive weighting strategies also appear in frameworks like PINNs [57] and Transformers [67], we show that the joint-minimization strategy is more effective in practice, primarily because the convex nature of the $\ell_{1}$ sparsity term over both $\bm{\Gamma}$ and $\bm{\Xi}$ precludes the existence of a saddle point required for the min-max formulation to converge reliably. As demonstrated in the sensitivity analysis (Sec. 3.7), both trainable hyperparameters $\bm{\Gamma}$ and $\lambda$ can be used for ADAM-SINDy optimization; however, prioritizing $\bm{\Gamma}$ is recommended, as it consistently recovers the correct terms with machine precision, whereas $\lambda$ may yield correct structures with slight numerical deviations. ADAM-SINDy supports both batch-wise and full-batch optimization. For simpler dynamical systems, small randomly sampled batches often capture sufficient temporal information to accurately estimate governing parameters. In contrast, complex systems may require larger batches to encompass the necessary information, making full-batch optimization preferable. Accordingly, full-batch training is adopted in this work, leading to reduced uncertainty in the converged parameters within the ADAM-SINDy framework.
Particular emphasis is placed on the essential steps required for the joint-minimization approach to succeed, given the counterintuitive nature of this optimization problem. First, $\bm{\Gamma}$ must be incorporated within the $\ell_{1}$ loss as a transformed value. In the present work, that transformation is the absolute value function, although alternatives such as an exponential functions are possible. Taking the absolute value of the $\bm{\Gamma}$ matrix prior to computing the loss forces $\bm{\Gamma}\to 0$ rather than -$\infty$. This absolute value function also introduces stochasticity into the trajectories of $\bm{\Gamma}$, as zero crossings occur, subsequently influencing $\bm{\Xi}$ and $\bm{\Lambda}$, despite the use of full-batch optimization in the examples. Second, sequential thresholding is required to select a sparse system within ADAM-SINDy. Without this step, all $\bm{\Gamma}$ values would tend toward zero, eventually reducing the problem to a least-squares fit after sufficient epochs. Thresholding allows for saturation of selected $\bm{\Gamma}$ values, thereby establishing a fixed minimum level of sparsity that can only increase with each epoch. The specific threshold value itself is an additional hyperparameter that may vary depending on the specific problem; however, the same threshold value was used across all examples in the present work, indicating robust behavior. The theoretical foundations of the joint-minimization approach are further elaborated in the B.
@@
Similarly, symbolic regression via neural networks has been employed to identify dynamical systems [60]. Boddupalli et al. [60] applied this approach to chemical reaction kinetics, the same system considered in this work (Sec. 3.4), resulting in an overly complex, “bloated” mathematical model when explicit stacking within the neural network is absent. However, when exponential terms are explicitly stacked with polynomials, the identified model closely approximates the ground truth equation but exhibits slight deviations in the coefficient values. Remarkably, ADAM-SINDy accurately identified the true dynamical system, capturing the dynamics without requiring additional treatments to ensure regression convergence.
@@
In real-world scenarios, observational data are inevitably contaminated by noise from various sources, motivating an evaluation of ADAM-SINDy’s efficacy under such conditions. To assess its robustness, we compared ADAM-SINDy with classical SINDy-W by introducing Gaussian white noise up to 50% of the standard deviation across all numerical examples. Although the results of these experiments are not included here, both methods exhibited comparable performance and failed to show clear advantages at high noise levels. At lower noise levels, both remained effective, suggesting that neither method consistently outperforms the other in the presence of significant noise. Future work will aim to enhance ADAM-SINDy’s resilience to noisy data.
A notable limitation of the full batch-ADAM-SINDy methodology is its high computational cost, stemming from the use of the ADAM optimizer for the coefficient matrix. While it offers advantages over classical SINDy, this improvement comes with considerable overhead, despite mirroring the computational strategies of its predecessor apart from the optimization algorithm. Efficiency can potentially be improved by employing an iterative least squares method for the coefficient matrix while retaining ADAM for nonlinear parameters and sparsity knobs, forming a combined least-squares–ADAM framework. However, integrating these methods poses challenges, particularly in aligning sparsity constraints with nonlinear optimization objectives. Further research is needed to develop robust methodologies that effectively merge these approaches while preserving their respective strengths. Notably, the ABC flow example demonstrated the benefit of embedding physical constraints directly into the learning process: incorporating a divergence-free condition into the loss function via the computational graph led to a dramatic reduction in training time from approximately 60,000 to just 40 epochs. This result highlights how leveraging underlying physical laws can not only enhance model fidelity but also significantly reduce computational cost, offering a promising direction for future work.
Given the primary focus on “concurrent nonlinear parameter computations” to address a key limitation of classical SINDy, ADAM-SINDy still does not fully overcome the challenge of constructing a master library containing all possible candidate terms for the numerical examples considered herein. Consequently, although a master library is included in the present problem setup, additional candidate functions were incorporated as needed for some specific problems. It is important to emphasize that this limitation is inherent to this class of problems. Classical SINDy acknowledged this constraint (see supplementary material in [34]) through an illustrative example involving sine series and polynomial summations. This example demonstrated that nonlinear functions tend to approximate higher-order polynomials as sparsity is enforced. Likewise, symbolic regression also begins the system identification with simple polynomial functions and proceeds through the evolution stage, progressively introducing higher-order polynomials and nonlinear terms. Thus, the development of new methodologies is necessary to overcome the difficulty of accurately identifying the correct candidate functions, despite the presence of high-order polynomials and complex nonlinearities.
@@ Acknowledgments
The ADAM-SINDy framework, developed in the PyTorch environment, is publicly available at https://github.com/siva-viknesh/ADAM-SINDy.
Appendix A ADAM-SINDy Optimization Algorithm
[…]
The parameters $\bm{\Xi}$, $\bm{\Lambda}$, $\lambda$, and $\bm{\Gamma}$ are then updated using stochastic gradient descent on a per-batch basis, progressively minimizing the loss function over the entire dataset. For full-batch optimization, updates are applied after evaluating the loss on the complete dataset. Importantly, the updates for the regularization parameters $\lambda$ and $\bm{\Gamma}$ differ from those of $\bm{\Xi}$ and $\bm{\Lambda}$. Depending on the optimization formulation, the regularization parameters are updated via either gradient ascent or gradient descent, governed by a sign variable $s$. Specifically, in the min-max optimization formulation of Eq. (8), $\lambda$ and $\bm{\Gamma}$ are updated through gradient ascent ($s=+1$), whereas in the joint minimization formulation of Eq. (9), the updates proceed via gradient descent (s=1s=-1). This flexible update methodology enables adaptive control of sparsity throughout training. After each batch update, a sequential thresholding step sets entries of both $\bm{\Xi}$ and $\bm{\Lambda}$ to zero if their magnitudes fall below a specified tolerance $\epsilon$. This iterative procedure continues for each batch until the maximum number of iterations $N_{\text{max}}$ is reached, ensuring efficient convergence to a sparse and accurate model representation.
Algorithm 1 ADAM-SINDy Algorithm
Input:
 Observational data $\mathbf{X}$ and $\dot{\mathbf{X}}$
 Tolerance $\epsilon$ and maximum iterations $N_{\text{max}}$
Initialization:
 Set iteration counter $n\leftarrow 0$
Optimization Loop:
while $n<N_{\text{max}}$ do
   For each batch $b=1,2,\dots$
    Sample a batch of data $\mathbf{X}_{b}$ and $\dot{\mathbf{X}}_{b}$
    Compute $\bm{\Theta}(\mathbf{X}_{b};\bm{\Lambda})$
    Compute Loss Function:
$\mathcal{L}_{b}\leftarrow\left\|\dot{\mathbf{X}}_{b}-\bm{\Theta}(\mathbf{X}_{b};\bm{\Lambda})\bm{\Xi}\right\|_{2}^{2}+\lambda\left\||\bm{\Gamma}|\bm{\Xi}\right\|_{1}$
    Update Parameters using Backpropagation:
    Update regularization parameters:
     where $s\leftarrow+1$ for ascent, $s\leftarrow-1$ for descent
       if Uniform Sparsity then
$\lambda_{n+1}\leftarrow\lambda_{n}+s\cdot\eta_{\lambda}\nabla_{\lambda}\mathcal{L}_{b}(\bm{\Theta}_{b},\bm{\Xi},\bm{\Lambda},\lambda,\bm{\Gamma})$
       else
$\bm{\Gamma}_{n+1}\leftarrow\bm{\Gamma}_{n}+s\cdot\eta_{\bm{\Gamma}}\nabla_{\bm{\Gamma}}\mathcal{L}_{b}(\bm{\Theta}_{b},\bm{\Xi},\bm{\Lambda},\lambda,\bm{\Gamma})$
   Increment iteration counter: $n\leftarrow n+1$
end while
Output:
 Optimized parameters $\bm{\Xi}^{*}$ and $\bm{\Lambda}^{*}$
Stochastic versus Full-Batch Optimization: The ADAM-SINDy framework provides the flexibility to implement both a batching strategy through stochastic gradient descent and a full-batch optimization approach. In the stochastic case, optimization is performed on smaller, randomly sampled subsets of the overall dataset. This method optimizes computational resource utilization by partitioning the data into manageable chunks, thereby reducing memory requirements and accelerating the optimization process. The batch size must be sufficiently large to capture the dynamics of interest while remaining small enough to facilitate efficient computation. In contrast, full-batch optimization processes the entire dataset in a single iteration. This approach leverages all available information, potentially yielding more accurate and stable updates to the optimization parameters. Notably, ADAM optimization employs momentum, which can help avoid local minima even in the absence of stochasticity introduced by mini-batches.
Appendix B Joint Minimization vs. Min-Max Optimization
In this section, we compare two optimization formulations considered within ADAM-SINDy: the min-max formulation and the joint minimization formulation. Both approaches aim to balance prediction accuracy and model sparsity but differ fundamentally in how sparsity is enforced. First, the min-max objective function is given by
$\min_{\mathbf{\Xi},\mathbf{\Lambda}}\max_{\mathbf{\Gamma}}\left[\left\|\dot{\mathbf{X}}-\mathbf{\Theta}(\mathbf{X};\mathbf{\Lambda})\mathbf{\Xi}\right\|_{2}^{2}+\lambda\left\||\mathbf{\Gamma}|\mathbf{\Xi}\right\|_{1}\right]\;,$
where the sparsity term is adversarially weighted by the variable $\mathbf{\Gamma}$. In contrast, the joint minimization formulation eliminates the adversarial structure by optimizing all variables cooperatively:
$\min_{\mathbf{\Xi},\mathbf{\Lambda},\mathbf{\Gamma}}\left[\left\|\dot{\mathbf{X}}-\mathbf{\Theta}(\mathbf{X};\mathbf{\Lambda})\mathbf{\Xi}\right\|_{2}^{2}+\lambda\left\||\mathbf{\Gamma}|\mathbf{\Xi}\right\|_{1}\right]\;,$
treating $\mathbf{\Gamma}$ as a passive regularization weight rather than an adversarial one. In both cases, the objective function consists of a data fidelity term (the $\ell_{2}$ norm) and a sparsity-promoting term (the $\ell_{1}$ norm). However, the role of $\mathbf{\Gamma}$ as either an adversary or a cooperative weight introduces fundamentally different optimization dynamics.
Theoretical analysis of the optimization frameworks
To analyze the optimization dynamics, we consider the gradients of the objective function with respect to $\mathbf{\Xi}$ and $\mathbf{\Gamma}$:
$\nabla_{\mathbf{\Xi}}L=-2\mathbf{\Theta}^{\top}(\dot{\mathbf{X}}-\mathbf{\Theta}\mathbf{\Xi})+\lambda\;|\mathbf{\Gamma}|\cdot\mathrm{sign}(\mathbf{\Xi}),$
$\nabla_{\mathbf{\Gamma}}L=\lambda\,|\mathbf{\Xi}|\cdot\mathrm{sign}(\mathbf{\Gamma}).$
where, $\mathrm{sign}(z)=\{+1\text{ if }z>0,\;0\text{ if }z=0,\;-1\text{ if }z<0\}$. These expressions highlight the fundamental instability in the optimization dynamics when $\mathbf{\Gamma}$ is treated adversarially (e.g., min-max). The gradient $\nabla_{\mathbf{\Gamma}}L$ drives $\mathbf{\Gamma}$ to grow unbounded in directions that increase the sparsity penalty $\||\mathbf{\Gamma}|\mathbf{\Xi}\|_{1}$, provided $\mathbf{\Xi}$ contains nonzero values. As $\mathbf{\Gamma}$ increases, the regularization term increasingly dominates $\nabla_{\mathbf{\Xi}}L$, subsequently driving $\mathbf{\Xi}\rightarrow 0$. This feedback loop leads to pathological dynamics in which $|\mathbf{\Gamma}|\rightarrow\infty$ and $\mathbf{\Xi}\rightarrow 0$, effectively suppressing model expressivity and hindering recovery of the correct terms. Importantly, this instability arises from the fact that the sparsity term is convex in both $\mathbf{\Xi}$ and $\mathbf{\Gamma}$. As a result, the overall objective lacks the convex-concave structure required for the existence of a well-defined saddle point. Specifically, a min-max problem admits a stable saddle point only when the objective is convex in the minimization variables and concave in the maximization variables. In our setting, the $\ell_{1}$ norm term is “convex” in both arguments $\mathbf{\Xi}$ and $\mathbf{\Gamma}$, rendering the adversarial maximization over $\mathbf{\Gamma}$ ill-posed. This contrasts with adversarial formulations successfully employed in other domains—such as self-adaptive PINNs [57]—where adaptive weighting is applied to loss terms that are ideally driven to zero (e.g., initial or boundary condition errors). In sparse regression, however, the $\ell_{1}$ penalty should not be driven to zero through increasing $\Gamma$ unless $\mathbf{\Xi}$ is identically zero, making adversarial maximization of the penalty inherently self-defeating. Without a natural saturation point, $\mathbf{\Gamma}$ continues to grow, leading to complete sparsity and failure to recover meaningful terms.
We do acknowledge that there are methods for artificially introducing a saddle point into the min-max loss landscape that can benefit the optimization. For instance, hyperparameters such as weight decay, can prevent excessive growth of $\Gamma$ leading to a maximum sparsity enforcement. Alternative optimization algorithms, as opposed to ADAM, such as optimistic gradient descent [68] or optimistic mirror descent [69], introduce negative gradient components that can also yield saddle points for optimization. However, in the sparse regression setting, these additional parameters proved to be sensitive and problem-specific through testing, limiting the robustness of the min-max formulation. Thus, even though the min-max optimization with these additional enhancements could solve these problems, we discourage its utilization.
In contrast, the joint minimization formulation avoids adversarial dynamics altogether, where the gradients $\nabla_{\mathbf{\Gamma}}L$ and $\nabla_{\mathbf{\Xi}}L$ act cooperatively to minimize the same objective. As a result, the gradient dynamics differ fundamentally. The parameter $\mathbf{\Gamma}$ no longer attempts to maximize the sparsity penalty but instead acts as a cooperative scaling factor on the $\ell_{1}$ regularization term. The minimization, in combination with the absolute value function on $\Gamma$, will always drive $\Gamma$ towards 0, differing from the unbounded growth in the min-max formulation. This movement towards zero is what allows for machine precision accuracy in nonlinear parameters as the impact of the $\ell_{1}$ penalty is reduced and as $\Gamma=0$, we reach a least-squares problem. We can also verify that $\Gamma$ will stabilize at nonzero values for incorrect candidate functions from the above $\nabla_{\mathbf{\Gamma}}L$ equation and the thresholding step, preventing complete removal of the $\ell_{1}$ penalty. When an incorrect candidate function is zeroed out due to sequential thresholding, the gradient becomes exactly zero, preventing further updates to $\nabla_{\mathbf{\Gamma}}L$. Therefore, $\mathbf{\Gamma}$ saturates at a nonzero value for the incorrect terms. However, the algorithm allows the $\Gamma$ values associated with correct candidate terms to continue moving towards 0, further increasing the relative impact of the $\ell_{2}$ loss, while preventing the discovered model sparsity from decreasing. It is additionally noteworthy that the trajectory of $|\bm{\Gamma}|$ does not decrease monotonically, contrary to intuition from the analytical gradient $\nabla_{\mathbf{\Gamma}}L$. Instead, it exhibits oscillations over the training epochs as the true sign of $\mathbf{\Gamma}$ flips from positive to negative, since achieving an exact zero value of $\mathbf{\Gamma}$ is practically impossible without explicit clipping. This oscillatory behavior also further justifies the incorporation of a scheduler (dynamic learning rate) to decrease the learning rate at specific epochs, with the resulting oscillations further scaled by minimum step size of $\eta_{\mathbf{\Gamma}}\cdot\epsilon$. For $\mathbf{\Gamma}$ to consistently move closer to zero and thus increase the $\ell_{2}$ impact on remaining model candidates, one of these parameters ($\epsilon$ or $\eta_{\mathbf{\Gamma}}$) must decrease. Reducing $\epsilon$ risks allowing incorrect terms to re-enter the model; therefore, it is preferable to reduce the learning rate $\eta_{\mathbf{\Gamma}}$ instead.
Although this joint minimization approach may appear mathematically ambiguous, it is consistent with the design of an adaptive weighting strategy—penalizing only incorrect terms while sparing correct ones. The sparsity parameters converge to zero for the correct terms (facilitating machine precision), whereas they saturate at a nonzero value for the incorrect terms. While this approach avoids the ill-posedness associated with adversarial optimization, it may reduce the efficacy of sparsity enforcement, particularly when the regularization parameter $\lambda$ is small or the data fidelity term dominates the objective. In such cases, the learned weights $\mathbf{\Gamma}$ may fail to sufficiently penalize small-magnitude components in $\mathbf{\Xi}$, resulting in denser models. Nevertheless, the joint formulation offers a more stable and tractable optimization landscape, making it better suited for robust and interpretable sparse regression in practice.
Empirical analysis of the optimization frameworks
Figure 8: Evolution of $\mathbf{|\Gamma|}$ over training epochs for all numerical examples considered in this work. Each plot shows the $\mathbf{|\Gamma|}$ trajectory for both randomly selected correct and incorrect terms, under both objective functions: joint minimization and min-max formulations.
To substantiate the theoretical discussion, we investigate the sparsity inducing $\bm{\Gamma}$ trajectory under both optimization formulations across all numerical examples considered in this work. Although the min-max formulation lacks a well-defined saddle point, we employ the optimistic gradient ascent/descent method—an algorithmic strategy developed to improve stability for adversarial training [70]. This method introduces a negative momentum term based on past gradients into the current updates, effectively mitigating oscillations and reducing the risk of divergence. Specifically, optimistic ADAM [68] is used for the maximization step, while vanilla ADAM is applied for the minimization step of the objective function.
A comparative analysis of the two objective formulations with respect to sparsity enforcement is presented in Fig. 8, where the evolution trajectories of $|\mathbf{\Gamma}|$ are plotted for two randomly selected candidate functions: one corresponding to a correct term and the other to an incorrect term. Across all experiments, the sparsity value of $\bm{\Gamma}$ associated with the incorrect term remains consistently higher than that of the correct term. This behavior aligns with the design of the adaptive weighting strategy, which aims to penalize incorrect candidates more strongly while applying minimal or no penalty to correct ones. Furthermore, despite identical initializations of $\bm{\Gamma}$, the min–max formulation imposes stronger sparsity constraints than the joint-minimization method, resulting in generally larger $\bm{\Gamma}$ values. Notably, this led to no candidate functions being selected under the min–max formulation, whereas the joint-minimization approach successfully identified correct terms as discussed in Sec. 3. We further observe that replacing optimistic ADAM with vanilla ADAM for the maximization step intensifies sparsity enforcement; however, these results are omitted for brevity.
While the min–max formulation may appear advantageous in theory, its lack of convex-concave structure results in unstable adversarial dynamics that hinder practical sparse regression. The joint-minimization framework, by avoiding these adversarial effects, provides a more stable and reliable optimization landscape that still enforces sparsity effectively. Thus, it serves as the preferred strategy within the ADAM-SINDy framework throughout this work.